Please answer the following problem: Use the point-slope form to write an equation of the line that passes through the point (5, -7) with a slope of 3/5.
Answer by stanbon(75887)

Use the point-slope form to write an equation of the line that passes through the point (5, -7) with a slope of 3/5.
---------------------
y+7 = (3/5)(x-5)
y = (3/5)x -3 -7
y = (3/5)x - 10
===================
Cheers,
Stan H.